Hi, after recalibration of TFW2 and fixing the slewing for TOFW, the DST are now being regenerated. The jobs will probably finish later today. There are a few runs to be aware of when one does PID on this data set: 13787-13788 (45 deg) and 13906 (40 deg) until end of run (13922), has the TFW2 at the short mrs distance. 13906 until end of run does not get new DSTs yet since the I've had problems with 2-3 peaks in the timeoffset calibration. This comes about when there is problem with the timing between trigger (according to Catalin). The problem even persist I do it with only trigger 3. Soon the AuAu_at_62 GTRs and DSTs will also be remade, as TFW2 has not been included in the data reduction so far. I'll keep you updated! Cheers, Truls -- *---------------------------------* |http://www.nbi.dk/~trulsml / |Truls Martin Larsen / |trulsml_at_nbi.dk . |The Niels Bohr Institute // |Work Address: / \0 |Blegdamsvej 17 /\_/ |DK-2100 Copenhagen / / |Tel: +45 353 25269 / -- | /_/ | |Home address: / \ |Holger Danskes Vej 28H| ' |DK-2000 Frederiksberg | |Denmark | |Mob: +45 20974802 | *----------------------* _______________________________________________ Brahms-dev-l mailing list Brahms-dev-l_at_lists.bnl.gov https://lists.bnl.gov/mailman/listinfo/brahms-dev-lReceived on Mon Jun 11 2007 - 06:41:32 EDT
This archive was generated by hypermail 2.2.0 : Mon Jun 11 2007 - 06:41:59 EDT